OpenStudy (anonymous):

Use completing of the square to find d:\[x^2-dx+5\]

OpenStudy (lgbasallote):

you want to turn this into a perfect square trinomial?

OpenStudy (anonymous):

Well, actually I just made that up... but I wanted to know how you would be able to find dx of a polynomial when given ax^2-bx+c, ax^2 and c

OpenStudy (lgbasallote):

oh nice question...you know that if you are given ax^2 + bx + c then (b/2)^2 = c right?

OpenStudy (anonymous):

Yes and thank you.

OpenStudy (lgbasallote):

so this time your b is d and c is 5 so \[\large (\frac{d}{2})^2 = 5\] \[\large \frac{d}{2} = \sqrt 5\] \[\large d = 2\sqrt 5\] got it?

OpenStudy (anonymous):

Yup...

OpenStudy (anonymous):

But can you do it on an hard example with integers or whole numbers...

OpenStudy (lgbasallote):

\[2x^2 - bx + 32\] try it :)

OpenStudy (anonymous):

\[4\sqrt{8}\]

OpenStudy (lgbasallote):

nope! :p you should get a whole number ;)

OpenStudy (anonymous):

b is still irrational :(

OpenStudy (lgbasallote):

remember in completing the square method it should be in the form x^2 + bx + c? that means you divide everything by the coefficient of x first

OpenStudy (anonymous):

hmm I see...

OpenStudy (lgbasallote):

tricky right ;)

OpenStudy (anonymous):

yes very... and I am stuck, haha

OpenStudy (lgbasallote):

you divide by 2 first that will give you x^2 - bx + 16 <--perfect square

OpenStudy (anonymous):

I don't divide b by 2?

OpenStudy (lgbasallote):

wait that should be x^2 - b/2 x +16 lol sorry

OpenStudy (anonymous):

Okay so b is 16

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

you got it!

OpenStudy (anonymous):

cool... Thanks @lgbasallote

OpenStudy (lgbasallote):

\[\Huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]

OpenStudy (anonymous):

Lol...[English accent] Why, Thank you fine sir...